OpenStudy (anonymous):

an equation for the line that contains (2,4) and has a undefined slope is : choices are: a. y=2x+4 b. x=2 c. y=3 d. y=0

OpenStudy (whpalmer4):

Any line with an undefined slope is a vertical line. A vertical line will be of the form \(x = k\) where \(k\) is a constant.

OpenStudy (anonymous):

I'm confused a little bit

OpenStudy (whpalmer4):

okay, ask me a question and I'll try to unconfuse you :-)

OpenStudy (whpalmer4):

first, do you understand why a vertical line has an undefined slope?

OpenStudy (anonymous):

yep

OpenStudy (whpalmer4):

okay, do you understand why a vertical line is of the form \(x = k\)?

OpenStudy (anonymous):

I think

OpenStudy (whpalmer4):

so what's the confusion? you've only got one answer choice which could possibly fit. y = 2x+4 is a line with slope 2 y = 3 is a horizontal line, slope 0 (can be written as y = 0x + 3) y = 0 is a horizontal line, slope 0 (can be written as y = 0x + 0) that leaves x = 2, which is a vertical line which passes through (2,4) as well as (2,5), (2,3), (2, 11152342523498234234), and other points :-)

OpenStudy (anonymous):

ooh ok thanks :D

OpenStudy (whpalmer4):

did that clear up the confusion?